OpenStudy (anonymous):

find domain of f(x)= \frac{\sqrt{1 + x}}{9 - x}

OpenStudy (anonymous):

\[f \left( x \right)=\frac{ \sqrt{1+x} }{ 9-x }\] for domain \[x+1\ge0,x \ge -1\] \[9-x \neq 0,9 \neq x~or x \neq 9\] domain is [-1,9) U (9, infinity )